Vinay:

Y=tan^-1(5ax/a^2/6x^2)

Ferredoxin4:

I'm sorry but dude you"re gonna need to write your equation a bit more clearer and what's the question.. Otherwise i bet it involves knowing the standard tan function..do you know what is the tan function and its features?

jhonyy9:

do you need simpliefy this or what ?

Gdeinward:

\[Y=\tan^{-1} (5ax \div a^{2}\div 6x^{2})\]

Gdeinward:

This I believe is the appropriate equation correct?

Gdeinward:

Nonetheless, we need more info. what do you think?

Gdeinward:

@Tranquility

Gdeinward:

the only thing you can do is simplify..... and that seems to be a very difficult thing to do.

Tranquility:

Agreed, we need more information on what exactly needs to be done. We also have an additional 'a' variable thrown in there which just makes it more perplexing

Gdeinward:

It is slightly different but might give us some insight.

Tranquility:

Searched it up https://questioncove.com/updates/51bd50cae4b0cfd1f31b3854 Here is the exact question \(\large \tan^{-1}(\frac{5ax}{a^2-6x^2})\) Show that \(\large \frac{dy}{dx}=\frac{3a}{a^2}+9x^2+\frac{2a}{a^2}+4x^2\)

Gdeinward:

The website requires membership. I will provide a screen shot.

Tranquility:

there could have been a typo and instead of a^2/6x^2 it was meant to be (a^2-6x^2)

Gdeinward:

makes perfect sense.

1 attachment
Gdeinward:

Thank you for your assistance @Tranquility